Suppose Qd=-37*P+866, Qs=34*P+10 for home and Qd*=-37*P+436, Qs*=34*P+34 for foreign. What is the free trade price to two decimal places?
Answer #1

Ascertaining equilibrium price at Home by equating demand and supply at home,

Qd = Qs

-37P + 866 = 34P + 10

-37P - 34P = 10 - 866

-71P = -856

P = 856/71 = 12.05

The equilibrium price at home is $12.05 per unit.

Ascertaining equilibrium price at Foriegn by equating demand and supply,

Qd = Qs

-37P + 436 = 34P + 34

-37P - 34P = 34 - 436

-71P = -402

P = 402/71 = 5.66

The equilibrium price in foriegn is $5.66 per unit.

Since, equilibrium price is lower in Foriegn country, in case of free trade, home will import and foriegn will export.

Calculate the import demand of home -

Import demand = Qd - Qs = -37P + 866 - (34P + 10) = -71P + 856

Calculate the export supply of foreign -

Export supply = Qs - Qd = 34P + 34 - (-37P + 436) = 71P - 402

Equating import demand and export supply to ascertain the free trade price

-71P + 856= 71P - 402

-71P - 71P = -402 - 856

-142P = -1258

P = 8.85

The free trade price is $8.85 per unit.